OpenStudy (anonymous):

Fan and Medal-------- Simplify the expression by combining like terms. -5(6j+7) - (2j+7)-4j+14

OpenStudy (anonymous):

First we expand the brackets -30j-35-2j-7-4j+14 Then we simplify = 22j-21

OpenStudy (anonymous):

Thats it?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

Lol thanks bro ur awesome , but on the answer sheet it shows -36j-28

OpenStudy (anonymous):

Careless mistake sorry I can't do my simple addition and subtraction lol -30j-35-2j-7-4j+14 = -36j-42+14 = -36j-28

OpenStudy (anonymous):

lmao its alright bro i appreciate it